# OmniPred: Language Models as Universal Regressors

### v4

@@ 1 Introduction
Regression is a fundamental task in experimental design, over many domains such as hyperparameter tuning, computer software, industrial engineering, and chemical discovery. The goal of regression is to predict a metric $y$ of a general system given a set of input features $x$. Such regressors can later be used for various applications, such as offline optimization (Kumar et al., 2022; Trabucco et al., 2022), online optimization (Cai et al., 2020), low-cost benchmarking (Zela et al., 2022; Eggensperger et al., 2015) and simulation (Mendis et al., 2019; Hashemi et al., 2018; Kaufman et al., 2021).
@@
Answering this question is highly important not only for the traditional field of experimental design, but also for the ever-changing field of LLM research, especially due to recent interest in the ability to forecast outcomes of complex systems (Gruver et al., 2023) and reward modeling in reinforcement learning fine-tuning (Ziegler et al., 2019). The textual processing abilities of LLMs are particularly attractive, as they can potentially bypass the need to tediously featurize inputs (i.e. the $x$’s) into raw numerical tensors. Prior to our work, there has been no such research specifically addressing the feasibility and utility of training a “universal” metric predictor over a large heterogenous offline dataset.
@@
  • To the best of our knowledge, we propose OmniPred, the first scalable yet simple metric prediction framework based on free-form textual representations, applicable to general input spaces.
@@ 2 Related Work and Motivation
Regressor Dynamic Input Spaces? Can Multitask? Tensorize? Rescale?
@@
These issues are summarized in Table 1. In principle, an ideal regressor should process $x$ and output $y$, both in absolute terms, independent of changing external statistics or search constraints. For example, if the underlying relationship is $y=x^{2}$, then the regressor’s prediction at $x=2$ should be invariant regardless if the constraint is $x\in[1,5]$ or $x\in[0,100]$. One way to accomplish this is to represent $x$ with an independent universal tokenizer without problem-specific numeric transformations (Zhou et al., 2023). This immediately unlocks a large amount of transferrability when dealing with variable-length inputs and additional contextual metadata.
Previous works in the token-based, or effectively text-to-text paradigm, have mostly been in reinforcement learning from human feedback (Ziegler et al., 2019), where regression over textual responses (the “$x$”), also known as reward modelling, has been crucial to the success of recent interactive LLMs such as ChatGPT (OpenAI, 2022) and Gemini (Google, 2024). While such works have shown that LLMs are able to imitate human ratings in the form of ordinal variables and their softmax probabilities (Bradley & Terry, 1952), they have not shown whether LLMs are capable of regression over precise numeric-based data where $y\in\mathbb{R}$.
@@ 3.2 Language Model
Sampling and Decoding: Through regular temperature decoding, we can repeatedly sample $\widehat{y}\sim s_{\theta}(x)$, to approximate the prediction distribution over $\mathbb{R}$. To remain robust to strong outliers, instead of using empirical mean, our $\text{Aggregate}(\cdot)$ function is defined to be the empirical median, since we found it leads to lower error from ablations in Section 6.1. Since the model may need to predict over unseen regions of the input space, we can also assess the model’s uncertainty by observing the concentration of sampled predictions $\widehat{y}$ and additionally specific log probabilities across every decoded token.
@@ 4.1 OSS Vizier Format
Task-level metadata $m$ consists of a title, owner username, description, objective name, and optional free-form text. Since the OSS Vizier API is meant to provide an optimization service for users, there can be many sources of transferrability due to user-specific settings. These include:
@@ 4.2 Datasets
Real World Data: To investigate metric prediction over real world data which contain a rich variety of tasks, we will use real world data from Google Vizier (Golovin et al., 2017). Because we are not limited to training on fully completed trajectories over flat input spaces, we can train on more data than just the 750K studies for training the OptFormer (Chen et al., 2022), as seen from Table 3.
@@ 5.2 Multi-task Transferrability
To verify whether the model is performing transfer learning by reading textual cues, in Table 4 we compare results against the case when data is “anonymized” using a study-dependent hash function. For BBOB, we hash metadata $m$ which originally displayed (function class, dimension, shift). For AutoML, we hash parameter names and string values. Each study can still be uniquely identified and trained over, but the model can no longer observe useful correlations from common textual clues. Interestingly, the model fails to train over the full anonymized BBOB dataset, a case when the data is too large and heterogeneous.
@@ 7 Discussion: Limitations and Extensions
In this work, our emphasis was to demonstrate the promise of applying language modelling to general-purpose regression, and thus our design choices remained relatively simple to avoid confounding factors. We list some limitations of our specific design, which opens many more potential areas of exploration.
In-Context Learning (ICL): By design, we maximized the allowed prompt length to allow flexibility in representing $(x,m)$ and thus did not use in-context regression, since a single $(x,m)$ prompt could be 10K+ tokens long for some applications. While online finetuning in principle allows infinite amounts of data to be absorbed at inference time, it is worth investigating ICL methods which allow arbitrarily long prompts as well. Current methods (Chen et al., 2022) require significant input compression are which only applicable to tabular-like $x$’s. Additional use of ChatGPT and other service-based chat APIs (Vacareanu et al., 2024) rely on regression as an emergent property of expensive LLM-based training, making their serious use difficult, especially as such services cannot absorb large amount of user-defined offline $(x,y)$ data which often contain the most relevant information for finetuning.
@@
Other Input Spaces: The OSS Vizier API primarily focuses on hyperparameter tuning spaces. Traditionally, more complex spaces such as combinatorics and graphs require sophisticated modeling techniques to form regressors, largely in part to difficulties in representing the $x$’s as tensors. In addition, many applications with non-expressible spaces such program synthesis are impossible to traditionally regress over. We believe that text and token-based representations are highly promising and are widely applicable to domains previously unexplored in the field of experimental design.
Other Metadata: While we performed ablations which anonymized $m$ and parameter names, more investigation could be made on what types of metadata are particularly useful for prediction. Such metadata could contain proxy metrics introduced by previous domain-specific works, such as Jacobian Covariance for neural architecture search (Mellor et al., 2021) and neural-network norms (Jiang et al., 2020) for the study of generalization. The relevant code implementing machine learning or programming tasks may be especially important.
@@ 8 Impact Statement
Our research does not involve such sensitive topics for regression, as it is performed over blackbox optimization data. We followed ethical guidelines as all users in our proprietary dataset have consented to have their tuning data saved and used for offline analysis. The proprietary real world dataset does not contain any sensitive personal information other than usernames. Furthermore, we do not released any of the trained checkpoints, as it may be possible to reverse-engineer parts of the training data, which can lead privacy violations and data leakage.
@@ B.2 Local Training
During local training, data is only sourced from a single study’s limited trials (at most 1000). The training set size can be lower than the batch size (256), and thus we must define one epoch as seeing the training data once, i.e. only one gradient step if training size $\leq$ batch size, but multiple gradient steps otherwise.
We use the same settings from pretraining for consistency, but allow a maximum of 30 epochs. For early stopping, validation loss is now measured over the entire validation set rather than sampled batches. Further specific changes:
@@ B.3 Inference
At inference time, we perform temperature sampling with a temperature of 1.0. We restrict the logits to only decode the custom floating point tokens for representing $y$-values. To maximize batch size for a 1x1 TPU V3, we generate 64 samples and select the empirical median of these floating point samples as our final prediction when computing error.
@@ C.2 Data Processing for Flat Space
A flat space is where every trial in the study specifies every parameter configured in the space. In this case, we convert the parameters into the unit hypercube $[0,1]^{d}$. For numeric parameters, we scale all values to $[0,1]$ range, using the linear scaling by default unless (reverse)-log scaling was configured. For CATEGORICAL parameters, we use a one-hot encoding.
@@ C.3 Data Processing for Conditional Space
A conditional space is where depending on one parameter’s value, another parameter may be unused by the blackbox function. Conditional spaces commonly appear in AutoML settings where different model classes require a different set of parameters to be tuned. Another common use case is when we wish to optimize a numeric hyperparameter in the log scale but include 0 in the search (e.g. dropout rate, regularization coefficient), i.e. $\{\texttt{UNUSED}\}\cup[l,u]$ where $l>0$.
@@ C.4 Regressor Baselines
We emphasize in our paper on the concept of flexible string-based regression using language models and large-scale multi-task training data, rather than claiming replacement over widely accepted baselines in single-task, apples-to-apples comparisons.
@@ D.1 Study Preprocessing
Removing bad trials: Users may ignore or fail to evaluate a proposed $x$. Furthermore, during some trials, the $y$-objective may be denoted with a special “infeasible” value (e.g. if a high batch size led to GPU out-of-memory, or training encountered NaNs). We remove such trials from consideration in our work, although one can extend our $y$-tokenization to support infeasibility in later works.

### v6

@@ 1 Introduction
Regression is a fundamental task for experimental design, in many domains such as hyperparameter tuning, computer software, industrial engineering, and chemical discovery. The goal of regression is to predict a metric $y$ of a general system given a set of input features $x$. Such regressors can later be used for various applications, such as offline optimization (Kumar et al., 2022; Trabucco et al., 2022), online optimization (Cai et al., 2020), low-cost benchmarking (Zela et al., 2022; Eggensperger et al., 2015) and simulation (Mendis et al., 2019; Hashemi et al., 2018; Kaufman et al., 2021).
@@
Answering this question is highly important not only for the traditional field of experimental design, but also for the ever-changing field of LLM research, especially due to recent interest in the ability to forecast outcomes of complex systems (Gruver et al., 2023) and reward modeling in reinforcement learning fine-tuning (Ziegler et al., 2019). The textual processing abilities of LLMs are particularly attractive, as they can potentially bypass the need to tediously featurize inputs (i.e. the $x$’s) into raw numerical tensors. Prior to our work, there has been no such research specifically addressing the feasibility and utility of training a “universal” metric predictor over a large heterogeneous offline dataset.
@@
  • We propose OmniPred, the first scalable yet simple metric prediction framework based on free-form textual representations, applicable to general input spaces.
@@ 2 Related Work and Motivation
Regressor Dynamic Input Spaces? Can Multitask? Tensorize $x$? Rescale $y$?
@@
These issues are summarized in Table 1. In principle, an ideal regressor should process $x$ and output $y$, both in absolute terms, independent of changing external statistics or search constraints. For example, if the underlying relationship is $y=x^{2}$, then the regressor’s prediction at $x=2$ should be the same regardless if the constraint is $x\in[1,5]$ or $x\in[0,100]$. One way to accomplish this is to represent $x$ with an independent universal tokenizer without problem-specific numeric transformations (Zhou et al., 2023). This immediately unlocks a large amount of transferrability when dealing with variable-length inputs and additional contextual metadata.
Previous works in the token-based, or effectively text-to-text paradigm, have mostly been in reinforcement learning from human feedback (Ziegler et al., 2019), where regression over textual responses (the “$x$”), also known as reward modeling, has been crucial to the success of recent interactive LLMs such as ChatGPT (OpenAI, 2022) and Gemini (Google, 2024). While such works have shown that LLMs are able to imitate human ratings in the form of ordinal variables and their softmax probabilities (Bradley & Terry, 1952), they have not shown whether LLMs are capable of regression over precise numeric-based data where $y\in\mathbb{R}$.
@@ 3.2 Language Model
Sampling and Decoding: Through regular temperature decoding, we can repeatedly sample $\widehat{y}\sim s_{\theta}(x)$, to approximate the prediction distribution over $\mathbb{R}$. To remain robust to strong outliers, instead of using empirical mean, our $\text{Aggregate}(\cdot)$ function is defined as the empirical median, since we found it leads to lower error from ablations in Section 6.1. Since the model may need to predict over unseen regions of the input space, we can also assess the model’s uncertainty by observing the concentration of sampled predictions $\widehat{y}$ and additionally specific log probabilities across every decoded token.
@@ 4.1 OSS Vizier Format
Task-level metadata $m$ consists of a title, username, description, objective name, and optional free-form text. Since the OSS Vizier API is meant to provide an optimization service for users, there can be many sources of transferrability due to user-specific settings. These include:
@@ 4.2 Datasets
Real World Data: To investigate metric prediction over real world data which contain a rich variety of tasks, we will use data collected by Google Vizier (Golovin et al., 2017). Because we are not limited to training on fully completed trajectories over flat input spaces, we can train on more data than just the 750K studies used in the OptFormer (Chen et al., 2022), as seen from Table 3.
@@ 5.2 Multi-task Transferrability
To verify whether the model is performing transfer learning by reading textual cues, in Table 4 we compare results against the case when data is “anonymized” using a study-dependent hash function. For BBOB, we hash metadata $m$ which originally displayed (function class, dimension, shift). For AutoML, we hash parameter names and string values. Each study can still be uniquely identified and trained on, but the model can no longer observe useful correlations from common textual clues. Interestingly, the model fails to train over the full anonymized BBOB dataset, a case when the data is too large and heterogeneous.
@@ 7 Discussion: Limitations and Extensions
In this work, our emphasis was to demonstrate the promise of applying language modeling to general-purpose regression, and thus our design choices remained relatively simple to avoid confounding factors. We list some limitations of our specific design, which opens many more potential areas of exploration.
In-Context Learning (ICL): By design, we maximized the allowed prompt length to allow flexibility in representing $(x,m)$ and thus did not use in-context regression, since a single $(x,m)$ prompt could be 10K+ tokens long for some applications. While online finetuning in principle allows infinite amounts of data to be absorbed at inference time, it is worth investigating ICL methods which allow arbitrarily long prompts as well. Current methods (Chen et al., 2022) which require significant input compression are only applicable to tabular-like $x$’s. Additional use of ChatGPT and other service-based chat APIs (Vacareanu et al., 2024) rely on regression as an emergent property of expensive LLM-based training, making their serious use difficult, especially as such services cannot absorb large amount of user-defined offline $(x,y)$ data which often contain the most relevant information for finetuning.
@@
Other Input Spaces: The OSS Vizier API primarily focuses on hyperparameter tuning spaces. Traditionally, more complex spaces such as combinatorics and graphs require sophisticated modeling techniques to form regressors, largely in part to difficulties in representing the $x$’s as tensors. In addition, many applications with non-expressible spaces such program synthesis are impossible to traditionally regress over. We believe that text and token-based representations are highly promising and widely applicable to domains previously unexplored in the field of experimental design.
Other Metadata: While we performed ablations which anonymized $m$ and parameter names, one can further investigate which types of metadata are particularly useful for prediction. Such metadata could contain proxy metrics introduced by previous domain-specific works, such as Jacobian Covariance for neural architecture search (Mellor et al., 2021) and neural-network norms (Jiang et al., 2020) for the study of generalization. The relevant code implementing machine learning or programming tasks may be especially important.
@@ 8 Impact Statement
Our research does not involve such sensitive topics for regression, as it is performed over blackbox optimization data. We followed ethical guidelines as all users in our proprietary dataset have consented to have their tuning data saved and used for offline analysis. The proprietary real world dataset does not contain any sensitive personal information other than usernames. Furthermore, we do not release any of the trained checkpoints, as it may be possible to reverse-engineer parts of the training data, which can lead to privacy violations and data leakage.
@@ 9 Conclusion
Acknowledgements
We would like to thank Olivier Bachem, Hado van Hasselt, John Jumper, Aviral Kumar, Yingjie Miao, Sebastian Nowozin, Mangpo Phothilimthana, Zi Wang, Scott Yak, and Amir Yazdanbakhsh for useful discussions and Daniel Golovin for continuing support.
@@ B.2 Local Training
During local training, data comes from a single study’s limited trials (at most 1000). The training set size can be lower than the batch size (256), and thus we must define one epoch as seeing the training data once, i.e. only one gradient step if training size $\leq$ batch size, but multiple gradient steps otherwise.
We use the same settings from pretraining for consistency, but allow a maximum of 30 epochs. For early stopping, validation loss is now measured over the entire validation set instead of sampled batches. Further specific changes:
@@ B.3 Inference
At inference time, we perform temperature sampling with a temperature of 1.0. We restrict the logits to only decode the custom floating point tokens for representing $y$-values. To maximize batch size for a 1x1 TPU V3, we generate 64 samples and select the empirical median of these floating point samples as our final prediction when computing prediction error.
@@ C.2 Data Processing for Flat Space
A flat space is where every trial in the study specifies every parameter configured in the space. In this case, we convert the parameters into the unit hypercube $[0,1]^{d}$. For numeric parameters, we scale all values into the range $[0,1]$, by default using linear scaling unless (reverse)-log scaling was configured. For CATEGORICAL parameters, we use a one-hot encoding.
@@ C.3 Data Processing for Conditional Space
A conditional space occurs when one parameter may be unused, depending on its parent parameter’s value. Conditional spaces commonly appear in AutoML settings where different model classes require a different set of parameters to be tuned. Another common use case is when we wish to optimize a numeric hyperparameter in the log scale but include 0 in the search (e.g. dropout rate, regularization coefficient), i.e. $\{\texttt{UNUSED}\}\cup[l,u]$ where $l>0$.
@@ C.4 Regressor Baselines
We emphasize that our paper’s contributions are mostly on regression using flexible string representations and large-scale multi-task training, and do not claim to replace widely accepted baselines in single-task, apples-to-apples comparisons.
@@ D.1 Study Preprocessing
Removing bad trials: Users may ignore or fail to evaluate a proposed $x$. Furthermore, during some trials, the $y$-objective may be denoted with a special “infeasible” value (e.g. if a high batch size led to GPU out-of-memory errors, or if training encountered NaNs). We remove such trials from consideration in our work, although one can extend our $y$-tokenization to support infeasibility in later works.